A falling body has an initial velocity of 200cm/seconds.How far will it drop before its velocity becomes 10000 cm/seconds.


1
Expert's answer
2021-01-29T13:26:42-0500

Well, we have given initial and final velocities and we have the value of g=9.81 m/s^2. We can find h (distance) easily:

"h=\frac{V^2(final)-V^2(initial)}{2\times g}" = "\frac{100^2-2^2}{2\times 9.81}=500.75 m"

It reaches to drop to 500.75 metr.